The legs of a right triangle are lengths x and x 3. The cosine of the smallest angle of the triangle is _____.
Found 2 solutions by Alan3354, GrantJones:
Answer by Alan3354(69443)

What does "x 3" mean?

Answer by GrantJones(3)

Now h = sqr(x^2 + x(sqr3)^2)
= sqr(x^2 + 3x^2)
= sqr(4x^2)
= 2x
Your answer is h = 2x